# Robust Multimodal Learning via Cross-Modal Proxy Tokens

### v2

@@ 4 Experiments and Results
Table 1: Performance comparison when modality gets missing during both training and inference. We train and evaluate our model with same modality ratio following [22, 16, 20] and report the mean and standard deviation (SD) of three independent runs with different seeds.
Datasets % availability ViLT MAP [22] MSP Visual- Ma et al. Kim et al. MuAP [6] CMPT (Ours)
Image Text [21] Input Attn. [16] BERT [23] [29] [20] Head Cross Mean $\pm$ SD
@@ 4.1 Datasets
We evaluate our approach on five popular multimodal datasets across different tasks. A brief description of each dataset is provided here, with further details in Section S2 in the supplementary.
UPMC Food-101 [46] is a multimodal classification dataset consisting of 101 food classes. It has a total of 90,704 image-text pairs, divided into a training set of 67,988 pairs and a test set of 22,716 pairs.
MM-IMDb dataset [2] contains 25,959 samples with image and text modalities for multi-label movie genre classification. It is split into 15,552 training, 2,608 validation, and 7,799 test samples across 23 classes.
Kinetics-Sound (KS) [1] is a subset of Kinetics400 dataset [17] for human action recognition using audio and video as input modalities. It has 31 classes, with 14,739 samples in the training set and 2,594 samples in the test set.
Audio-Visual Event (AVE) localization dataset [42] is used for multimodal event localization. It includes 4,143 10-second videos divided into train/val/test splits containing 3,339/402/402 videos, respectively. The dataset spans 28 event categories.
CREMA-D dataset [4] includes 7,442 short video clips from 91 actors expressing six emotion categories. It contains 6,698 samples for training and 744 for testing. The dataset is used for multimodal emotion recognition.
@@ 4.2 Implementation details
Vision-Language datasets. We use pretrained ViT-B [8] from CLIP model [38] as image encoder and BERT-base-uncased model [27] as text encoder. The max token length is set to 40 for UPMC Food-101 and 256 for MM-IMDb dataset, respectively.
Audio-Video datasets. We use AST model [13] pretrained on AudioSet dataset [12] as audio encoder and ViT-B [8] from CLIP model [38] as video encoder. Raw audio is converted to a 128-bin mel spectrogram at 16 kHz, with a maximum length of 1024 frames. For video, we sample 3 frames randomly while training and uniformly while testing. Other pre-processing steps are similar to [55, 36, 9].
Hyperparameter settings. We use Python 3.8.19 and PyTorch 2.2.2 for training and evaluating our models. All the models are trained using two NVIDIA RTX 2080Ti GPUs. For vision-language datasets, we set the learning rate to $10^{-3}$ and train the models for 10 epochs with a batch size of 8. For audio-video datasets, the learning rate is set to $5\times 10^{-5}$ and models are trained for 100 epochs with a batch size of 4. We use AdamW [26] optimizer with $\epsilon=10^{-8}$ and weight decay = 0.02. Cross entropy loss is applied while training the models. We utilize a polynomial learning rate scheduler with power=0.9 and set the first 5 epoch as warm-up. The learning rate is set to 0.1 times the original learning rate during warm-up epochs. We set LoRA [15] rank = 1 and insert them after query, key, value and output layers of each transformer block. Further details with all the hyperparameters can be found in Section S3 in the supplementary.
@@
Table 3: Performance comparison with both modalities available during training and inference. We report % accuracy for UPMC Food-101 and F1-Macro/F1-Micro score for MM-IMDb dataset. ‘-’ indicate results that are not reported in corresponding papers.
Methods Parameters (M) UPMC MM-IMDb
@@
ViLT [21] 112.63 112.63 92.00 55.30 / 64.70
MBT [33] 196.00 196.00 93.56 59.60 / 64.81
MMBT [19] 170.00 170.00 94.10 60.80 / 66.10
MMLoRA [9] 196.10 196.10 93.70 61.70 / 67.20
PMF [24] 198.43 2.54 91.51 58.77 / 64.51
PMF-L [24] 643.95 4.44 91.68 61.66 / 66.72
MLA [55] 218.26 218.26 93.33 - / -
@@ 4.3.1 Missing modality during training and inference
In the first setup, a modality is missing at a fixed rate during both training and inference, as studied in [22, 16, 20]. We follow the experiment setup in [22, 16, 20] and compare our method against seven state-of-the-art approaches on the MM-IMDb and UPMC Food-101 datasets. As shown in Table 1, our method consistently outperforms all baselines across various missing modality configurations. Specifically, compared to the most recent approach in Kim et al. [20], we achieve a 4.18%–9.40% improvement in F1-macro score on MM-IMDb dataset and a 0.16%–1.56% improvement in accuracy on UPMC Food-101 dataset across different missing modality scenarios. These results also suggest that the CMPTs can provide good proxy for the missing class tokens, which significanctly improves the model robustness. To ensure a reliable evaluation, we run our method with three different random seeds and report the mean and standard deviation (SD) for various missing scenarios. The low SD values indicate the stability and reliability of our method. We also assessed the performance of the models with different distributions of modality ratios at training and testing, for which the results are presented in Section S4 of the supplementary material.
@@ 4.3.2 Missing modality during inference only
In the second setup, all modalities are available during training, but one modality is completely missing during inference, as examined in [55, 39]. Following the setup in [55], we evaluate our model when all modalities are available during training, but one modality is completely missing during inference. As shown in Table 2, our method outperforms existing baselines in most cases. For instance, when the audio modality is missing, all methods except MLA [55] experience a significant performance drop on the CREMA-D and Kinetics-Sound (KS) datasets. On the CREMA-D dataset, our method achieves 7.93% and 11.30% improvements over the most recent MLA approach, when video and audio are missing, respectively. Similarly, on the KS dataset, our method outperforms MLA with 13.60% and 34.74% improvements when video and audio are missing, respectively. While MLA performs better when the image modality is missing on the UPMC Food-101 dataset, it suffers a larger performance drop when text is missing. In contrast, our method maintains a balanced performance across different missing modalities. Furthermore, across all datasets, our method achieves the best performance when all modalities are available and demonstrates superior robustness compared to other approaches. These results highlight the effectiveness of our method in handling scenarios where a modality is entirely absent during inference.
Table 4: Performance (accuracy) comparison with both modalities available during training and inference on Audio-Video datasets. denotes results generated using existing code.
Method Parameters (M) KS AVE CREMA-D
Total Learnable
G-Blend [45] 22.36 22.36 62.20 65.50 58.70
FiLM [37] 22.35 22.35 63.33 - 60.07
BiGated [18] 22.35 22.35 63.72 - 59.21
OGM-GE [36] 22.35 22.35 65.74 76.90 68.14
QMF [54] 22.36 22.36 65.78  68.16 63.71
PMR [11] 22.36 22.36 - 66.40 61.80
MLA [55] 22.37 22.37 71.35  64.68 79.70
MBT [19] 172.00 172.00 85.00 77.80 -
MMLoRA [9] 172.19 172.19 91.40 96.20 88.60
CMPT (Ours) 172.16 0.17 91.21 96.77 88.84
(a) Dataset: MM-IMDb
@@
Figure 3: Generalization to varying missing rates during inference. Models are trained with 100% image + 100% text and evaluated with 100% image + $x$text following [29, 20]. Our approach demonstrates better generalization, particularly under severe modality loss.
@@ 4.4 Performance when all modalities are available
A method designed to handle missing modalities should not compromise performance when all modalities are present. In this subsection, we demonstrate that our approach also achieves comparable or superior performance to state-of-the-art models across five datasets when all modalities are available. These results highlights that incorporating CMPTs does not introduce any drawback when all modalities remain available.
To evaluate this, we train and test our method when all modalities are available during training and inference. Results in Table 3 demonstrate that our proposed method achieves state-of-the-art performance on both the UPMC Food-101 and MM-IMDb datasets. Our method surpasses the most recent MLA approach [55] by 1.14% on UPMC Food-101. Our method achieves 1.88% and 2.03% improvement in F1-macro and F1-micro scores, respectively, over second-best performing MMLoRA [9] on MM-IMDb.
We present the results for audio-video datasets in Table 4, where we compare our method against nine state-of-the-art methods. Our approach consistently achieves comparable or superior performance across all three datasets, while using the fewest number of learnable parameters.
4.5 Generalization to different missing rates
In real-world scenarios, the missing rates for modalities during inference can be unpredictable. Any method designed to handle missing modalities needs to generalize effectively across varying levels of missing modality rates during inference. To evaluate this capability, we trained our models on complete multimodal data and assessed their performance under varying levels of missing modalities during inference. Specifically, we follow the experiment setup similar to [29, 20] where models are trained with 100% image + 100% text data and evaluated with 100% image + $x$text.
As illustrated in Figure 3, our method consistently outperforms all baselines across different missing scenarios, demonstrating superior robustness. We observe a decline in performance of all methods as the amount of available text decreases, but our approach maintains stronger resilience compared to other methods. For instance, on the MM-IMDb dataset, our method achieves a 5.8% higher F1-Macro score than the most recent method by Kim et al. [20] and a 25.6% improvement over ViLT [21] when only 10% of text is available. A similar trend is observed on the UPMC Food-101 dataset, where our method consistently yields superior performance, confirming its robustness to missing modalities at inference.
Note that published work in [29, 20, 21] did not report results for different missing rates of images, which prevents us from reporting a direct comparison in that setting. Nevertheless, our results highlight the effectiveness of CMPTs in handling different missing modality scenarios, making it a more generalizable solution for real-world applications.
@@ 4.5 Generalization to different missing rates
Figure 4: Effectiveness of CMPT in handling missing modalities on MM-IMDb dataset. CMPT significantly improves performance when modalities are severely missing. All models are trained on 100% image + 100% text and tested with varying levels of missing modalities.
@@
Figure 5: CMPT improves performance across most of the classes when modalities are missing. However, it fails to accurately predict certain classes (e.g., short/documentary) in the absence of text, likely due to the image modality’s insufficient information for those classes. Classes are sorted based on the amount of performance improvement.
@@ 4.6.1 Effectiveness of CMPT
To evaluate the effectiveness of CMPTs in handling missing modalities, we compare our method against two baselines: (i) a standard model trained without modality dropout or CMPT (baseline) and (ii) a model trained with modality dropout, where modalities are randomly removed during training to improve robustness to missing data.
@@
Figure 4 presents the main results for MM-IMDb dataset, where the baseline model experiences a sharp performance drop as the proportion of missing modalities increases. While modality dropout improves robustness, it fails to fully recover the lost performance. Incorporating CMPT further enhances performance, achieving a 32.06% improvement over the baseline and 11.55% over modality dropout when the image modality is entirely missing as shown in Figure 4(a). A similar trend is observed when text is missing as shown in Figure 4(b). Notably, CMPTs become increasingly effective as the amount of missing data increases, demonstrating its strong ability to mitigate performance degradation in the presence of severe modality loss. Similar trend appears in UPMC Food-101 dataset, which we discuss in Sec. S5 in the supplementary material.
4.6.2 Effects of CMPT on different classes
@@
When analyzing the effectiveness of CMPT, a key question is whether the overall performance gains are driven by just a few classes or if CMPT consistently improves performance across most classes. To address this, we conducted a per-class performance analysis with and without CMPT.
@@
As shown in Figure 5, CMPT improves performance across nearly all classes in both text-missing and image-missing scenarios. Notably, it significantly mitigates performance degradation in modality-sensitive classes. For example, when text is missing, classes like Horror, Animation, and War experience severe drops in performance, but CMPT effectively recovers much of the loss (Figure 5(a)). Similarly, when images are missing, CMPT provides substantial gains in Thriller, Action, and Romance (Figure 5(b)). A slight drop is observed in Short and Documentary classes when text is missing, likely due to insufficient information in the image modality alone for those classes. These results demonstrate that CMPT enhances performance across most classes, leading to a strong overall performance improvement. A similar trend is observed on the UPMC Food-101 dataset (further discussed in Section S6 of the supplementary material).

### v4

@@ 4 Experiments and Results
Table 1: Performance comparison when modality gets missing during both training and inference. We train and evaluate our model with same modality ratio following existing baselines (Lee et al., 2023; Jang et al., 2024; Kim & Kim, 2024) and report the mean and standard deviation (SD) of three independent runs with different seeds.
Datasets % availability ViLT MAP (2023) MSP VisualBERT Ma et al. Kim & Kim MuAP (2024) CMPT (Ours)
Image Text (2021) Input Attn. (2024) (2019) (2022) (2024) Head Cross Mean $\pm$ SD
@@ 4.1 Datasets
We evaluate our approach on five popular multimodal datasets across different tasks. A brief description of each dataset is provided here, with further details in Section A2 in the appendix.
UPMC Food-101 (Wang et al., 2015) is a multimodal classification dataset consisting of 101 food classes. It contains 90,704 image-text pairs, divided into a training set of 67,988 pairs and a test set of 22,716 pairs.
MM-IMDb (Arevalo et al., 2017) has 25,959 samples with image and text modalities for multi-label movie genre classification. It is split into 15,552 training, 2,608 validation, and 7,799 test samples across 23 classes.
Kinetics-Sound (KS) (Arandjelovic & Zisserman, 2017) is a subset of Kinetics400 dataset (Kay et al., 2017) for human action recognition using audio and video as input modalities. It has 31 classes, with 14,739 samples in the training set and 2,594 samples in the test set.
Audio-Visual Event (AVE) dataset (Tian et al., 2018) contains 4,143 10-second videos for multimodal event localization. It is divided into train/val/test sets containing 3,339/402/402 samples across 28 categories.
CREMA-D dataset (Cao et al., 2014) includes 7,442 short video clips from 91 actors expressing six emotion categories. It contains 6,698 training and 744 test samples for multimodal emotion recognition.
@@ 4.2 Implementation details
Vision-Language datasets. We use pretrained ViT-B (Dosovitskiy et al., 2021) from CLIP model (Radford et al., 2021) as image encoder and BERT-base-uncased model (Lu et al., 2019) as text encoder. The max token length is set to 40 for UPMC Food-101 and 256 for MM-IMDb dataset, respectively.
Audio-Video datasets. We use AST model (Gong et al., 2021) pretrained on AudioSet dataset (Gemmeke et al., 2017) as audio encoder and ViT-B (Dosovitskiy et al., 2021) from CLIP model (Radford et al., 2021) as video encoder. Raw audio is converted to a 128-bin mel spectrogram at 16 kHz, with a maximum length of 1024 frames. For video, we sample 3 frames randomly while training and uniformly while testing. Other pre-processing steps are similar to (Zhang et al., 2024; Peng et al., 2022; Du et al., 2023).
Hyperparameter settings. We use Python 3.8.19 and PyTorch 2.2.2 for training and evaluating our models. All the models are trained using two NVIDIA RTX 2080Ti GPUs. For vision-language datasets, we set the learning rate to $10^{-3}$ and train the models for 10 epochs with a batch size of 8. For audio-video datasets, the learning rate is set to $5\times 10^{-5}$ and models are trained for 100 epochs with a batch size of 4. We use AdamW (Loshchilov & Hutter, 2019) optimizer with $\epsilon=10^{-8}$ and weight decay = 0.02. While training, we utilize cross entropy loss, polynomial learning rate scheduler with power=0.9 and set the first 5 epoch as warm-up. The learning rate is set to 0.1 times the original learning rate during warm-up epochs. We set LoRA (Hu et al., 2022) rank = 1 and insert them after query, key, value and output layers of each transformer block. Further details with all the hyperparameters can be found in Section A3 in the appendix.
@@
Table 3: Performance comparison with both modalities available during training and inference on Image-Text datasets. We report % accuracy for UPMC Food-101 and F1-Macro/F1-Micro score for MM-IMDb dataset. Here, ‘-’ indicate results that are not reported in corresponding papers.
Methods Parameters (M) UPMC $M$-IMDb
@@
ViLT (2021) 112.63 112.63 92.00 55.30 / 64.70
MBT (2021b) 196.00 196.00 93.56 59.60 / 64.81
MMBT (2019) 170.00 170.00 94.10 60.80 / 66.10
MMLoRA (2023) 196.10 196.10 93.70 61.70 / 67.20
PMF (2023) 198.43 2.54 91.51 58.77 / 64.51
PMF-L (2023) 643.95 4.44 91.68 61.66 / 66.72
MLA (2024) 218.26 218.26 93.33 - / -
@@
Table 4: Performance (% accuracy) comparison with both modalities available during training and inference on Audio-Video datasets. Here, ‘’ denotes results generated using existing code.
Method Parameters (M) KS AVE CREMA-D
Total Learnable
G-Blend (2020a) 22.36 22.36 62.20 65.50 58.70
FiLM (2018) 22.35 22.35 63.33 - 60.07
BiGated (2018) 22.35 22.35 63.72 - 59.21
OGM-GE (2022) 22.35 22.35 65.74 76.90 68.14
QMF (2023b) 22.36 22.36 65.78  68.16 63.71
PMR (2023b) 22.36 22.36 - 66.40 61.80
MLA (2024) 22.37 22.37 71.35  64.68 79.70
MBT (2019) 172.00 172.00 85.00 77.80 -
MMLoRA (2023) 172.19 172.19 91.40 96.20 88.60
CMPT (Ours) 172.16 0.17 91.21 96.77 88.84
@@ 4.3.1 Missing modality during training and inference
In the first setup, a modality is missing at a fixed rate during both training and inference, as studied by Lee et al. (2023); Jang et al. (2024); Kim & Kim (2024). We follow their experiment setup and compare our method against seven state-of-the-art approaches on MM-IMDb and UPMC Food-101 datasets. As shown in Table 1, our method consistently outperforms all baselines across various missing modality configurations. Specifically, compared to the most recent approach by Kim & Kim (2024), we achieve a 4.18%–9.40% improvement in F1-macro score on MM-IMDb dataset and a 0.16%–1.56% improvement in accuracy on UPMC Food-101 dataset across different missing modality scenarios. It is worth noting that while the performance of Kim & Kim (2024) on UPMC Food-101 dataset is comparable to ours in two scenarios, our method still performs best in every configuration. These results also suggest that the CMPTs can provide good proxy for the missing class tokens, which significantly improves the model robustness. To ensure a reliable evaluation, we run our method with three different random seeds and report the mean and standard deviation (SD) for various missing scenarios. The low SD values indicate the stability and reliability of our method. We also assessed the performance of the models with different distributions of modality ratios at training and testing, for which the results are presented in Section A5 of the appendix.
@@ 4.3.2 Missing modality during inference only
In the second setup, following the work of Zhang et al. (2024), we evaluate our model when all modalities are available during training, but one modality is completely missing during inference. As shown in Table 2, our method outperforms existing baselines in most cases. For instance, when the audio modality is missing, all methods except MLA (Zhang et al., 2024) experience a significant performance drop on the CREMA-D and Kinetics-Sound (KS) datasets. On the CREMA-D dataset, our method achieves 7.93% and 11.30% improvements over the most recent MLA approach, when video and audio are missing, respectively. Similarly, on the KS dataset, our method outperforms MLA with 13.60% and 34.74% improvements when video and audio are missing, respectively. While MLA performs slightly better when the image modality is missing on the UPMC Food-101 dataset, it suffers a larger performance drop when text is missing. In contrast, our method maintains a balanced performance across different missing modalities. Furthermore, across all datasets, our method achieves the best performance when all modalities are available and demonstrates superior robustness compared to other approaches. These results highlight the effectiveness of our method in handling scenarios where a modality is entirely absent during inference.
(a) Dataset: $M$-IMDb
@@
Figure 2: Generalization to varying missing rates during inference. Models are trained with 100% image + 100% text and evaluated with 100% image + xx% text following Ma et al. (2022); Kim & Kim (2024). Our approach demonstrates better generalization, particularly under severe modality loss.
@@ 4.4 Performance when all modalities are available
A method designed to handle missing modalities should not compromise performance when all modalities are present. In this subsection, we demonstrate that our approach also achieves comparable or superior performance to state-of-the-art models across five datasets when all modalities are available.
To evaluate this, we train and test our method when all modalities are available during training and inference. Results in Table 4 demonstrate that our proposed method achieves state-of-the-art performance on both UPMC Food-101 and MM-IMDb datasets. Our method surpasses the most recent MLA approach (Zhang et al., 2024) by 1.14% on UPMC Food-101. Our method achieves 1.88% and 2.03% improvement in F1-macro and F1-micro scores, respectively, over second-best performing MMLoRA (Du et al., 2023) on MM-IMDb.
We present the results for audio-video datasets in Table 4, where we compare our method against nine state-of-the-art methods. Our approach consistently achieves comparable or superior performance across all three datasets, while using the fewest number of learnable parameters. These results highlights that incorporating CMPTs does not introduce any drawback when all modalities remain available.
@@
Figure 3: Effectiveness of CMPTs on the MM-IMDb dataset. All models are trained with 100% image and 100% text data, then evaluated under varying amount of missing modalities. CMPTs achieve significant performance improvement, especially under severe missing modality scenarios.
@@
Figure 4: CMPTs improve performance across most of the classes when modalities are missing. However, it fails to accurately predict certain classes (e.g., short/documentary) in the absence of text, likely due to the image modality’s insufficient information for those classes. Classes are sorted based on the amount of performance improvement.
4.5 Generalization to different missing rates
In real-world scenarios, the missing rates for modalities during inference can be unpredictable. Any method designed to handle missing modalities needs to generalize effectively across varying levels of missing modality rates during inference. To evaluate this capability, we trained our models on complete multimodal data and assessed their performance under varying levels of missing modalities during inference. Specifically, we follow the experiment setup similar to Ma et al. (2022); Kim & Kim (2024) where models are trained with 100% image + 100% text data and evaluated with 100% image + xx% text.
As illustrated in Figure 2, our method consistently outperforms all baselines across different missing scenarios, demonstrating superior robustness. We observe a decline in performance of all methods as the amount of available text decreases, but our approach maintains stronger resilience compared to other methods. For instance, on the MM-IMDb dataset, our method achieves a 5.8% higher F1-Macro score than the most recent method by Kim & Kim (2024) and a 25.6% improvement over ViLT (Kim et al., 2021) when only 10% of text is available. A similar trend is observed on the UPMC Food-101 dataset, where our method consistently yields superior performance, confirming its robustness to missing modalities at inference.
Note that published work by Ma et al. (2022); Kim & Kim (2024); Kim et al. (2021) did not report results for different missing rates of images, which prevents us from reporting a direct comparison in that setting. Nevertheless, our results highlight the effectiveness of CMPTs in handling different missing modality scenarios, making it a more generalizable solution for real-world applications.
@@ 4.6.1 Effectiveness of CMPTs
To evaluate the effectiveness of CMPTs in handling missing modalities, we compare our method against two baselines: (i) a standard model trained without modality dropout or CMPTs (baseline) and (ii) a model trained with modality dropout, where modalities are randomly removed during training to improve robustness to missing data. For a fair comparison, all models were fine-tuned on the downstream task using LoRA with a rank of 1.
@@
Figure 3 presents the main results for MM-IMDb dataset, where the baseline model experiences a sharp performance drop as the proportion of missing modalities increases. While modality dropout improves robustness, the performance gap remains large. Incorporating CMPTs further enhance performance, achieving a 32.06% improvement over the baseline and 11.55% over modality dropout when the image modality is entirely missing as shown in Figure 3(a). A similar trend is observed when text is missing as shown in Figure 3(b). Notably, CMPTs become increasingly effective as the amount of missing data increases, demonstrating its strong ability to substantially mitigate performance degradation in the presence of severe modality loss. Similar trend appears in UPMC Food-101 dataset, which we discuss in Sec. A6 in the appendix.
4.6.2 Effects of CMPTs on different classes
@@
When analyzing the effectiveness of CMPTs, a key question is whether the overall performance gains are driven by just a few classes or if CMPTs consistently improves performance across most classes. To address this, we conducted a per-class performance analysis with and without CMPTs.
@@
As shown in Figure 4, CMPTs improve performance across nearly all classes in both text-missing and image-missing scenarios. Notably, it significantly mitigates performance degradation in modality-sensitive classes. For example, when text is missing, classes like Horror, Animation, and War experience severe drops in performance, but CMPTs effectively recover much of the loss (Figure 4(a)). Similarly, when images are missing, CMPTs provide substantial gains in Thriller, Action, and Romance (Figure 4(b)). A slight drop is observed in Short and Documentary classes when text is missing, likely due to insufficient information in the image modality alone for those classes. These results demonstrate that CMPTs enhance performance across most classes, leading to a strong overall performance improvement. A similar trend is observed on the UPMC Food-101 dataset which is discussed in Section A7 of the appendix.
Table 5: Ablation study on the alignment loss weight $\lambda$. We report performance on the UPMC Food-101 and AVE datasets across both complete and missing modality scenarios.
Dataset Available Modality Missing Modality $\lambda=0.0$ $\lambda=0.1$ $\lambda=0.2$ $\lambda=0.3$ $\lambda=0.4$ $\lambda=0.5$
UPMC Food-101 Image Text 73.12 73.48 75.66 73.72 74.39 74.46
Text Image 84.31 84.79 85.31 84.51 84.81 84.88
Image - Text - 93.96 94.14 94.47 94.13 94.34 94.21
AVE Audio Video 84.82 84.33 85.21 84.82 84.33 84.82
Video Audio 82.83 83.08 84.58 85.32 81.84 84.83
Audio - Video - 95.77 96.02 96.77 96.52 96.52 96.27
4.6.3 Determining the Optimal Alignment Loss Weight $\lambda$
To determine the optimal value for the alignment loss weight $\lambda$, we performed an ablation study on the UPMC Food-101 and AVE datasets, varying $\lambda$ from 0.0 to 0.5. The results, shown in Table 5, indicate that the choice of $\lambda$ is crucial for performance in different scenarios. For our experiments, $\lambda=0.20$ provides an overall better performance compared to other values. On the UPMC Food-101 dataset, setting $\lambda=0.20$ consistently achieves the best performance across all conditions: with missing text, missing image, and with both modalities available. Using $\lambda=0.0$, which is equivalent to removing the alignment loss, results in a significant performance drop, highlighting the importance of the alignment objective.
On the AVE dataset, the results show a similar trend. A weight of $\lambda=0.20$ yields the best performance when video is missing and when both modalities are available. $\lambda=0.30$ shows a slight performance improvement in the video-only scenario. Given that $\lambda=0.20$ demonstrates superior or near-optimal performance consistently across both datasets, we set $\lambda=0.20$ for all our experiments, as it provides a balance in performance across different modality conditions.
Figure 5: t-SNE plots of the fused feature tokens $\mathcal{T}$ for ten classes under different modality settings. The t-SNE plots show that fused features without CMPTs (red) deviate significantly from the complete multimodal features (green). Incorporating CMPTs (blue) provides embeddings that closely align with the complete multimodal features, indicating improved semantic alignment and robustness.
4.6.4 Qualitative Analysis of CMPT Effectiveness
To gain deeper insights into how CMPTs enhance robustness in missing modality scenarios, we conducted a qualitative analysis on the UPMC Food-101 dataset using both fused feature projections and attention visualizations. Figure 5 presents t-SNE plots of the fused feature $\mathcal{T}$ for ten representative food classes where CMPTs yield notable performance gains. We compare three settings: (i) complete multimodal input (green), (ii) missing modality without CMPTs (red), and (iii) missing modality with CMPTs (blue). The t-SNE plots show that fused features without CMPTs (red) deviate significantly from the complete multimodal features (green) in missing modality scenarios. Incorporating CMPTs (blue) provides embeddings that closely align with the complete multimodal features, indicating improved semantic alignment and robustness.
We also analyze the attention maps from the last layer of the vision encoder for the CLS token and CMPTs in the text-missing scenario, as shown in Figure 6. We selected random examples from six classes where fused features without CMPTs provide incorrect predictions, but including the CMPTs leads to correct predictions (approximately 15% samples in the dataset show this behavior). The attention maps show that CLS tokens attend to semantically-irrelevant regions, which likely leads to incorrect predictions. In contrast, CMPTs attend to semantically-relevant regions for the respective classes and lead to accurate predictions. These visualizations demonstrate that CMPTs not only act as proxies for the missing modality but also reshape the model’s attention to reflect modality-relevant cues. For the sake of completeness, we have also included attention maps for some cases when the fused features with and without CMPTs provide correct predictions (approximately 60% samples show this pattern) and when the fused features with CMPTs lead to incorrect predictions (approximately 3% samples show this pattern) in Section A8.
Figure 6: Attention map visualizations from the last layer of the vision encoder for the CLS token and CMPTs in the text-missing scenario on the UPMC Food-101 dataset. The attention maps show that CLS tokens attend to semantically-irrelevant regions, which likely leads to incorrect predictions. In contrast, CMPTs attend to semantically-relevant regions for the respective classes and lead to accurate predictions.